Chapter 06, Problem 53 Starting from rest, a 98-kg firefighter slides down a fire pole. The average frictional force evertedi need help ASAP
Answer #1

From work energy theorem we have, w-=[5mw7+mgh,) - 6 mo + mgh) -sch = {m} = m(o)-65m(0) + msha) Ich = = mv,?-mghe mgh – Suh =